A 0.635-g sample of cobalt, Co, is added to 235 mL of 0.100 M H2SO4, and reacts according to

Co(s)+H2SO4(aq)⟶CoSO4(aq)+H2(g)

What is the limiting reactant?

H2SO4(aq) OR Co(s)

A volume of 155 mL of hydrogen is collected over water; the water level in the collecting vessel is the same as the outside level. Atmospheric pressure is 756.0 Torr and the temperature is 25 °C.

What is the percent yield of hydrogen for this reaction?

percent yield:
Answer #1

The balanced equation

Balanced equation:
Co(s) + H2SO4(aq) ===> CoSO4(aq) + H2(g)

Reaction type: single replacement

Mass of Cobalt = 0.635 gm

Molar mass of cobalt = 58.933 gm /mol

Moles of Cobalt = 0.635 gm / 58.933 gm /mol = 0.01077 Moles

Moles of H2SO4 = 235 x 0.1 / 1000 = 0.0235 Moles

Limiting reagent is Cobalt . Because it is exits in less moles  (0.01077 only)

Theoretical Moles of H2 produced = 0.01077 Moles

Actual moles of H2 produced

PV= nRT

P = Pressure in atm                       V= Volume in Liter

n = no of moles               R = 0.0821 L atm K-1 Mol-1

T = Temperature in Kelvin

P = 756 torr = 0.9947 atm V= 155 ml n = ? T = 398 K

n = PV / RT = 0.9947 x 155 / (0.08206 x 398 ) = 4.72 Moles

Moles of H2 actually produced (4.72 moles) is more than theoretical value (0.01077 Moles) and percentage yield is more than 100 %. It is not possible at all. Please recheck the given values.